cotx lim (4 (4x+1) x-oton Ot (4 kt t il
Answer #1

Cotx 6 lim (4x+13) Julma tigas) = CA - X-) ot sol By lemit chain rule 1 g(x) = cofa do 14x+1), f(a)=en d dx Incirax) da bocusplease rate me

Thanks